what is the relationship between two goods if marginal rate of substitution between them is zero



1
Expert's answer
2021-02-22T17:31:24-0500

Goods that have its marginal rate of substitution between them being zero are termed complementary goods. To find the marginal rate, the marginal utility of product 1 is divided by the marginal utility of product 2 (Becchetti et al., 2020). The outcome should be the same as the change in quantity of product 2 divides it by a change in the quantity of product 1.

MRS1,2 =MU1/MU2= Δ quantity of product 2/Δg quantity of product 1

The indifference curve will be in the form of L shape. The Marginal Rate of substitution of vertical indifference curve will be 

The indifference curve (IC) of complementary good is L shape. 

MRS of IC (Vertical) = Δ quantity of product 2/0=0

or 

0/Δquantity of product 1= 0


Reference

Becchetti, L., Bruni, L., & Zamagni, S. (2020). The microeconomics of wellbeing and sustainability: Recasting the economic process.